OpenStudy (anonymous):

Okay, guys new one :) If a right triangle has a side length of 11 and a hypotenuse length of 12, how long is the third side? 11^2 + 12^2=c^2 121+144=√265?

OpenStudy (anonymous):

Yay or nay?

OpenStudy (lgbasallote):

nope...hypotenuse should be c

OpenStudy (lgbasallote):

so it's 11^2 + b^2 = 12^2

OpenStudy (anonymous):

Oh, yeah...okay! Give me a sec

OpenStudy (anonymous):

√23?

Parth (parthkohli):

Yepyepyep

OpenStudy (lgbasallote):

144 - 121 = 23 so you're right!

OpenStudy (anonymous):

:D Thanks! :D

Parth (parthkohli):

yw

Parth (parthkohli):

I didn't even do anything lol why am I even welcoming

OpenStudy (anonymous):

haha well you said yep....

OpenStudy (lgbasallote):

no...he said yepyepyep...makes the difference

OpenStudy (anonymous):

haha *true story*

Parth (parthkohli):

lgba surpassed my SS